In triangle ABC, the measure of angle B is 21 degrees less than the measure of angle A multiplied by 4. The measure of angle C is 1 degree more then the measure of angle A multiplied by 5. Determine the measure of each interior angle and each exterior angle of triangle ABC.
Answer by fcabanski(1391)

B=4A-21
C=5A+1
A=A


The above is given. The interior angles of a triangle add to 180.


4A-21 + 5A+1 + A =180


10A-20=180


10A = 200


A = 20


B=4A-21=4(20)-21=59


C=5A+1=5(20)+1=101


When you know the measure of all interior angles, you can find the two exterior angels adjacent to each interior angle by subtracting the interior angle from 180.


The two exterior angles adjacent to A = 180-20=160


The two exterior angles adjacent to B = 180=59=121


The two exterior angles adjacent to C = 180-101=79

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com